In the given op amp circuit, R1 = 3 R and R2 = 4 R. Determine the voltage gain vo/vs. Assume that R = 10 kΩ.

The voltage gain vo/vs is
Answer #1

Ib 30kn 40kh W B Ia R B R WWW Toka 10K © + of OV .+ VaL and R=4R R,=3R - 3X1OKR R1 = 30kr & R=40kr For ideal op-amp any It-I- 3 Vs = Vs-ke II) Apply .: Vx=LIVS nodal ar analysis certing terminal at inver of epamp2 Vx-o (10k) o-vo (406) uve ol 9 ( -4